Grant is five times as old as Francis. One third of the sum of their ages is equal to 32. Find their ages.
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

Grant is five times as old as Francis. One third of the sum of their ages is equal to 32. Find their ages.
..
Francis = x years
Grant = 5x
sum=6x
..
(1/3)*6x=32
multiply by 3
6x=32*3
/6
x= 32*3/6
x=16 Francis's age
Grant's age =80
..
m.ananth@hotmail.ca

Answer by stanbon(75887)

Grant is five times as old as Francis. One third of the sum of their ages is equal to 32. Find their ages.
-----------------------------------
Equations:
g = 5f
(1/3)(g+f) = 32
---
Rearrange:
g=5f
g+f = 96
----
Substitute for "g" and and solve for "f":
5f + f = 96
6f = 96
f = 16 (Francis' age now)
---
Since g = 5f, g = 5*16 = 80 (Grant's age now)
================================================
Cheers,
Stan H.